A cars covers a distance of 2 km in 2.5 minutes . If it covers half of the distance with speed 40 km/HR the rest distance it shall covers with what speed

Grade:11

2 Answers

riya
6 years ago
2 km= 2.5 min = 150 s
 
half distance i,e  1 km speed is 40 km/hr
                            for  1km it takes  60*60/40 = 90 s
            total time taken = 150 s
           for first half distance = 90 s
      then therest will be   150 s – 90 s =60 s 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
bhumika
6 years ago
given;
distance=2.km
time=2.5minutes=0.041 hr
also given
half of distance ii.e 1km is travelled with speed 40 km/hr
wkt
speed=distance/time
time=distance/speed
=1/40=0.025
..nowtime taken to complete rest of the journey =0.041-0.025=0.016 hr
now,
speed=1/0.016=62.5km/hr ~ 60 km/hr